# Accommodating Fabrication Defects on Floquet Codes with Minimal Hardware Requirements

### v1

@@ 5.1 Accommodating defective qubits on a square lattice
It is possible to implement Floquet codes such as the honeycomb code on a square lattice, through only using a subset of available connections on the device. An example of this is given in Fig. 9(a), with unused connections shown in white. We are also therefore able to accommodate defective qubits without these additional connections using the strategy described in Sec. 3, as shown in Fig. 9(b). As before, this techique involves reducing the shrink plaquettes to weight-2 and combine four merge plaquettes into one super-plaquette.
However, there is nothing stopping us from using these additional connections in the quantum hardware. In Fig. 9(c), we instead reduce the shrink plaquettes to weight-4, and combine only two merge plaquettes together. This can be done by using two additional connections available on the square lattice. The result is a smaller super-plaquette, leading to a higher distance: the horizontal distance is reduced by one and the vertical distance is unchanged, compared to both being reduced by one when not using extra connectivity.
@@
Figure 9: Removing a defective qubit from a Floquet code implemented on a square lattice. (a) An implementation of the honeycomb code on a square lattice. White edges denote edges which are not used in the experiment. A defective qubit is shown in red. (b) Removing the defective qubit using the algorithm described in Sec. 3. This method uses the same connections as used by the original honeycomb code. (c) Removing the defective qubit using two edges which are not used in the honeycomb code but are available in the square lattice to reduce the size of the super-plaquette.
@@ 5.2 Accommodating defective qubits on a heavy hex lattice
Alternatively, we are able to treat the defective auxiliary qubit as a defective connection. In Appendix G, we present an approach for removing defective connections from Floquet codes while maintaining the same properties of tri-valence and 3-face-colourability. By using this approach to remove just the defective connection rather than the data qubits incident to that connection, we are able to create the Floquet code shown in Fig. 10(c). This approach removes no data qubits, and only removes three auxiliary qubits. It also only requires one plaquette to be split into three weight-2 plaquettes, and only three plaquettes need to be merged into a super-plaquette. It reduces both horizontal and vertical distances by one, whereas the code shown in Fig. 10(b) has vertical distance reduced by two.
@@ 6 Conclusion
While we have discussed how to implement this strategy when there are hardware defects known at the start of the experiment, one area that needs further work is accommodating defects which occur in the middle of an experiment, such as via high-energy radiation, qubit loss or erasure noise. Qubit loss in the middle of the experiment, are common in both photonic and neutral atom systems [5, 6, 43, 8]. Different proposals for correcting erasure events (which can also model qubit loss) in Floquet codes have been studied, through either modifying the stabilisers around the erased qubits [44], or through introducing additional Pauli error mechanisms [45]. A comparison of these methods to our code modifications around defects would be useful for Floquet code implementation in hardware. Aasen et al. also have proposals for identifying defective qubits mid-experiment [22]. We believe our work can also be adapted to accommodate defects which occur mid-experiment, but leave this for future work.
@@ Appendix A Improvements with extra connectivity
With the above strategy, a single defect reduces either the horizontal or vertical Pauli distance by one, while leaving the other unchanged. Moreover, we can accommodate the defect while removing five connections from the code lattice, and adding two, resulting in a net loss of three connections. We show in Appendix B that any strategy for removing a defective qubit must removal at least three connections, meaning that this approach is optimal in this regard. Given these improvements, we therefore expect that the performance of Floquet codes with defective qubits, given this extra connectivity, would be improved: the percolation threshold would likely be higher, and the performance under Pauli noise likely better.
@@ Appendix B Super-plaquette algorithm and minimal removal of qubits, plaquettes and edges
We begin this Appendix by commenting on why the algorithm presented in Sec. 3.1 preserves the important features of the lattice for defining the updated Floquet code: tri-valence and 3-face-colourability. We do this for the bulk of the code and note that the result follows for the boundary by cutting out the patch used, as explained in the main text. The first three steps of the algorithm simply involve selecting shrink and merge plaquettes. There will be exactly two shrink plaquettes. Let $K_{1}$ (resp. $K_{2}$) be the weight of the first (resp. second) shrink plaquette. The first (resp. second) shrink plaquette is adjacent to $K_{1}/2$ (resp. $K_{2}/2$) merge plaquettes, with two merge plaquettes being incident to both shrink plaquettes. Therefore the total number of merge plaquettes is $(K_{1}+K_{2})/2-2$.
@@
Now notice that the edges in shrink plaquettes of type $T\neq D$ (where $D$ is the defect edge basis) do not feature in any plaquette after the algorithm is applied: they are removed from the lattice (see Fig. 2). However, every qubit adjacent to such an edge in the original lattice (apart from the two removed qubits) is adjacent to an additional edge introduced in Step 4 of the algorithm. Hence, each vertex in the updated lattice remains adjacent to three edges as required.
@@
Now, for 3-face-colourability, note that the two-body plaquettes are guaranteed to be adjacent to the super-plaquette, call it $S_{1}$, and one other plaquette from the original lattice, call it $S_{2}$. By 3-face-colourability of the original lattice, since a two-body plaquette inherits the colouring of the parent shrink plaquette, it is coloured differently to $S_{1}$ (which by construction is the same colour as a merge plaquette) and to $S_{2}$. Meanwhile, the super-plaquette is adjacent to two-body plaquettes (which we have seen must be coloured differently to it) or to plaquettes of the original lattice adjacent to a merge plaquette, which we assume are coloured differently to it. Hence both types of newly added plaquette are adjacent to no other plaquettes of the same colour in the updated lattice. This is sufficient to conclude that the resulting lattice remains 3-face-colourable.
We will now show that the algorithm in Sec. 3.1 is optimal in terms of qubits, plaquettes and edges removed. As before, we will focus on the issue of removing a defective qubit in the bulk of the code. For the purpose of this proof, we will assume that code has periodic boundary conditions, meaning that the underlying graph is on a torus. Note that a code with planar boundary conditions can be embedded within a larger code with periodic boundary conditions [3]. This means that an optimal strategy for accommodating defects in the bulk of a planar Floquet code cannot perform better than an optimal strategy for accommodating defects in a Floquet code with periodic boundary conditions.
@@
Tri-valence on the torus requires that $E=3V/2$, where $E$ is the number of edges and $V$ is the number of vertices in the graph. This constraint is incompatible with an odd number of vertices, and hence we cannot remove just a single qubit. If two qubits are removed, the number of vertices becomes $\tilde{V}=V-2$, and tri-valence dictates that the number of edges becomes $\tilde{E}=3\tilde{V}/2=3V/2-3$.
@@
The Euler characteristic on a torus is $\chi=V-E+F=0$, where $F$ is the number of plaquettes. Having removed two qubits and three edges, the Euler characteristic becomes
@@
meaning that in order to maintain $\tilde{\chi}=0$, we must have $\tilde{F}=F-1$. In summary, a scheme for removing a defective vertex from the lattice while maintaining 3-face-colourability and tri-valence must remove at least two vertices, three edges and one face. Therefore, the algorithm outlined in Sec. 3.1 is optimal in terms of qubit, edge and plaquette removal.
@@
In our scheme, we remove certain edges from the honeycomb lattice and add extra edges along existing connections in the hardware, meaning that we effectively remove some connections from the lattice (see white highlighted edges in Fig. 2(c)). This is necessary if we do not rely on extra connectivity in the hardware (i.e., we do not have edges between qubits that are not connected in the honeycomb lattice). Indeed, our scheme is also optimal in terms of the connections that are removed from the lattice.
@@
Figure 12: Vertices which violate the tri-valence requirements of Floquet codes when removing a defective qubit. (a) After removing edges incident to the defective qubit and its neighbour, four non-adjacent vertices in green are degree-2, thus requiring additional edges. (b) After adding more edges to correct the degree-2 vertices, four vertices in blue are degree-4, meaning that edges incident to these vertices must be removed.
@@
To see this, we first note that we must remove edges that are adjacent to removed qubits (there are five of these). An example is shown in Fig. 12(a). Note that we now have four vertices of degree two, which are coloured in green. To maintain tri-valence after this removal, we must include at least four extra edges, since there are four degree-2 vertices in the resulting lattice, and none of these vertices are adjacent to each other. Indeed these extra edges must lie along existing connections in the lattice in order to avoid requiring additional connections in hardware. Adding these extra edges makes these vertices degree-3, but now leads to four vertices of degree four. One instance is shown in Fig. 12(b), with the degree-4 vertices in blue.
@@
In order to reduce these vertices to degree-3, we need to remove at least two edges from the original graph, as each edge removed will reduce the degree of the vertices it is incident to by one [^1] Note that the construction of Floquet codes used in this paper does not allow for self-loops, so removing an edge cannot reduce the degree of a vertex by two.. Hence, our scheme is also optimal in terms of the connections that are removed from the lattice. Note that we do not claim optimality for non-trivial combinations of defects, since then one must, for example, carefully choose the defect edges to achieve the optimal result. We do not have an optimal algorithm for such a choice, rather we use simplifying heursitics towards optimality (see Appendix D).
@@ Appendix C Applicability to other Floquet codes
In this appendix, we discuss how the techniques described in Sec. 3 can be applied to other 2-D Floquet codes. Many 2-D Floquet codes can be defined on a colour code lattice, including the 4.8.8 Floquet code, the hyperbolic Floquet codes and the CSS Floquet code [4, 48, 49, 24, 23]. These codes can be adapted to accommodate defective qubits by selecting a defect edge incident to the defective qubit, along with selecting the merge and shrink plaquettes in the same way as in Steps 2 and 3 as described in Sec. 3.1, with the rest of the super-plaquette algorithm following naturally. The only step that may change is Step 4, wherein we choose bases for the added edges. However, if one can make an identification of edge basis with edge colour (see, e.g., Ref. [23]), then this modification is simple.

### v2

@@ 5.1 Accommodating defective qubits on a square lattice
It is possible to implement Floquet codes such as the honeycomb code on a square lattice, through only using a subset of available connections on the device. An example of this is given in Fig. 9(a), with unused connections shown in white. We are also therefore able to accommodate defective qubits without these additional connections using the strategy described in Sec. 3, as shown in Fig. 9(b). As before, this techique involves reducing the shrink plaquettes to weight-2 and combining four merge plaquettes into one super-plaquette.
However, we can improve on this strategy by using these additional connections in the quantum hardware. In Fig. 9(c), we instead reduce the shrink plaquettes to weight-4, and combine only two merge plaquettes together. This can be done by using two additional connections available on the square lattice. The result is a smaller super-plaquette, leading to a higher distance: the horizontal distance is reduced by one and the vertical distance is unchanged, compared to both being reduced by one when not using extra connectivity.
@@
Figure 9: Removing a defective qubit from a Floquet code implemented on a square lattice. (a) An implementation of the honeycomb code on a square lattice. White lines denote connections which are not used in the experiment. A defective qubit is shown in red. (b) Removing the defective qubit using the algorithm described in Sec. 3. This method uses the same connections as used by the original honeycomb code. (c) Removing the defective qubit using two edges which are not used in the honeycomb code but are available in the square lattice to reduce the size of the super-plaquette.
@@ 5.2 Accommodating defective qubits on a heavy hex lattice
Alternatively, we are able to treat the defective auxiliary qubit as a defective connection. In Appendix G, we present an approach for removing defective connections from Floquet codes while maintaining the same properties of tri-valence and 3-face-colourability. Our strategy works by selecting one plaquette incident to the defective connection to be the shrink plaquette, and the other to be a merge plaquette. Other plaquettes which are adjacent to the shrink plaquette and of the same colour as the merge plaquette are also selected to be merge plaquettes. With the shrink plaquette and merge plaquettes chosen, Steps 4-7 of the algorithm in Sec. 3.1 can then be followed, with $D$ being the basis of the defective edge. In Appendix G, we prove the following:
Theorem 3 (for defective auxiliary qubits).
The above process produces a lattice such that (a) each data qubit is incident to three auxiliary qubits, (b) each auxiliary qubit is incident to two data qubits, and (c) the lattice is 3-face-colourable.
By using this approach to remove just the defective connection rather than the data qubits incident to that connection, we are able to create the Floquet code shown in Fig. 10(c). This approach removes no data qubits, and only removes two additional auxiliary qubits alongside the defective one. In Appendix G, we show that this is also optimal for uniform heavy lattices without requiring additional connectivity:
Theorem 4 (for defective auxiliary qubits).
Consider a defective auxiliary qubit in the bulk of a Floquet code defined on a heavy hex, heavy 4.8.8 or heavy 4.6.12 lattice (without boundary); i.e., a uniform tiling of the plane. Under the constraint that we cannot form an edge between any two qubits not connected by an edge in the original lattice, the algorithm described above removes the defective auxiliary qubit with the minimal number of removed auxiliary qubits from the original lattice.
@@ 6 Conclusion
While we have discussed how to implement this strategy when there are hardware defects known at the start of the experiment, one area that needs further work is accommodating defects which occur in the middle of an experiment, such as via high-energy radiation, qubit loss or erasure noise. Qubit loss in the middle of the experiment is common in both photonic and neutral atom systems [6, 7, 43, 9]. Different proposals for correcting erasure events (which can also model qubit loss) in Floquet codes have been studied, through either modifying the stabilisers around the erased qubits [44], or through introducing additional Pauli error mechanisms [45]. A comparison of these methods to our code modifications around defects would be useful for Floquet code implementation in hardware. Aasen et al. also have proposals for identifying defective qubits mid-experiment [23]. We believe our work can also be adapted to accommodate defects which occur mid-experiment, but leave this for future work.
@@ Appendix A Improvements with extra connectivity
With the above strategy, a single defect reduces either the horizontal or vertical Pauli distance by one, while leaving the other unchanged. Moreover, we can accommodate the defect while removing five connections from the code lattice, and adding two, resulting in a net loss of three connections. We show in Appendix B that any strategy for removing a defective qubit must remove at least three edges, meaning that this approach is optimal in this regard. Given these improvements, we therefore expect that the performance of Floquet codes with defective qubits, given this extra connectivity, would be improved: the percolation threshold would likely be higher, and the performance under Pauli noise likely better.
@@ Appendix B Super-plaquette algorithm and minimal removal of qubits, plaquettes and edges
In this Appendix we prove certain features of the algorithm presented in Sec. 3.1. We begin by commenting on why the algorithm preserves the important features of the lattice for defining the updated Floquet code: tri-valence and 3-face-colourability. We do this for the bulk of the code and note that the tri-valence and 3-face-colourability result follows for the boundary by cutting out the patch used, as explained in the main text. We also prove optimality, in terms of removed qubits, edges and plaquettes, in the bulk. There may be subtleties associated with optimality at the boundary (e.g., one could redefine the boundary around the defect), which we do not deal with here.
Proof of Theorem 1.
@@
The first three steps of the algorithm simply involve selecting the defect edge, and the shrink and merge plaquettes. There will be exactly two shrink plaquettes§§§If $K_{1}$ (resp. $K_{2}$) is the weight of the first (resp. second) shrink plaquette, then the total number of merge plaquettes is $(K_{1}+K_{2})/2-2$. The first (resp. second) shrink plaquette is adjacent to $K_{1}/2$ (resp. $K_{2}/2$) merge plaquettes, with two merge plaquettes being incident to both shrink plaquettes.. Now notice that the edges in shrink plaquettes of type $T\neq D$ (where $D$ is the defect edge basis) do not feature in any plaquette after the algorithm is applied: they are removed from the lattice (see Fig. 2). However, every qubit incident to such an edge in the original lattice (apart from the two removed qubits) is incident to an additional edge introduced in Step 4 of the algorithm. Hence, each vertex in the updated lattice remains incident to three edges as required.
@@
Now, for 3-face-colourability, note that the two-body plaquettes are guaranteed to be adjacent to the super-plaquette, call it $S_{1}$, and one other plaquette from the original lattice, call it $S_{2}$. By 3-face-colourability of the original lattice, since a two-body plaquette inherits the colouring of the parent shrink plaquette, it is coloured differently to $S_{1}$ (which, by construction, is the same colour as a merge plaquette—the merge plaquettes are coloured differently to both shrink plaquettes) and to $S_{2}$ (which was also adjacent to the parent shrink plaquette). Meanwhile, the super-plaquette is adjacent to two-body plaquettes (which we have seen must be coloured differently to it) and to plaquettes of the original lattice adjacent to a merge plaquette, which we assume are coloured differently to it. Hence both types of newly added plaquette are adjacent to no other plaquettes of the same colour in the updated lattice. This is sufficient to conclude that the resulting lattice remains 3-face-colourable.
@@
We will now show that the algorithm in Sec. 3.1 is optimal in terms of qubits, plaquettes and edges removed. We assume (as above) that the lattice has no boundaries. Let us it to be on a torus—this can be changed without changing the result. Since our scheme maintains tri-valence and 3-face-colourability, the number of removed plaquettes and edges are constrained by the number of removed qubits. Tri-valence on the torus requires that $E=3V/2$, where $E$ is the number of edges and $V$ is the number of vertices in the graph. This constraint is incompatible with an odd number of vertices, and hence we cannot remove just a single qubit. Our algorithm is therefore optimal in terms of the number of removed qubits, removing as it does two qubits. The rest follows by necessity. If two qubits are removed, the number of vertices becomes $\tilde{V}=V-2$, and tri-valence dictates that the number of edges becomes $\tilde{E}=3\tilde{V}/2=3V/2-3$. Removing more vertices results in more edges being removed: hence our algorithm is optimal in terms of the number of removed edges as well.
@@
If the removal of the defect does not change the number of logical qubits then the Euler characteristic of the graph must not change. The Euler characteristic on a torus is $\chi=V-E+F=0$, where $F$ is the number of plaquettes. Having removed two qubits and three edges, the Euler characteristic becomes
@@
meaning that in order to maintain $\tilde{\chi}=0$, we must have $\tilde{F}=F-1$. Once again, removing more vertices requires the removal of more faces. In summary, a scheme for removing a defective vertex from the lattice while maintaining 3-face-colourability and tri-valence must remove at least two vertices, three edges and one face. Therefore, the algorithm outlined in Sec. 3.1 is optimal in terms of qubit, edge and plaquette removal. ∎
@@
In our scheme, we remove certain edges from the honeycomb lattice and add extra edges along existing connections in the hardware, meaning that we effectively remove some connections from the lattice (see white highlighted edges in Fig. 2(c)). This is necessary if we do not rely on extra connectivity in the hardware (i.e., we do not have edges between qubits that are not connected in the honeycomb lattice). Indeed, our scheme is also optimal in terms of the connections that are removed from the lattice, at least when considering a single defect removed from the honeycomb, 4.8.8 or 4.6.12 lattices (the tri-valent, 3-face-colourable lattices based on a uniform tiling of the plane by regular polygons).
@@
Figure 12: Vertices which violate the tri-valence requirements of Floquet codes when removing a defective qubit. (a) After removing edges incident to the defective qubit and its neighbour, four non-adjacent vertices are degree-2, thus requiring additional edges. These vertices are coloured green. These are connected in pairs by shortest paths in the lattice (coloured dark green and light blue respectively). (b) After adding more edges to correct the degree-2 vertices, four vertices in blue are degree-4, meaning that edges incident to these vertices must be removed.
Proof of Theorem 2.
@@
From the above, we must remove two qubits per defect in the lattice. We first note that we must remove edges that are adjacent to these two removed qubits. An example is shown in Fig. 12(a). Note that we now have four vertices of degree two (coloured green in Fig. 12(a)). There are four of these because there are exactly five removed edges (one of which is the defect edge) and each pair of qubits is connected by at most one edge in the original lattice. Call the set of these degree-2 vertices $S$. To maintain tri-valence after this edge removal, we must include extra edges incident to these vertices in $S$. We assume the edges in the original lattice, before the defect removal, account for all possible available connections in the device. Hence, the extra edges we add must lie along existing connections in the lattice in order to avoid requiring additional connections in hardware.
In order to avoid introducing any vertices of degree-4 upon adding extra edges (see the blue vertices in Fig. 12(b)), we need to remove some more edges from the graph. The way this must work is as follows. Firstly, without loss of generality, let the defect edge be coloured red. Then two of the vertices in $S$ were adjacent to a green removed edge and two were adjacent to a blue removed edge (see Fig. 12(a); the former lie on green plaquettes in the figure and the latter on blue plaquettes). Let these pairs of vertices belong to subsets $S_{G}$ and $S_{B}$ respectively. Notice that the vertices in $S_{G}$ ($S_{B}$) are not incident to any green (blue) edge (see Fig. 12(a)). Thus we must add more green (blue) edges to the graph. To do so while maintaining tri-valence and 3-face-colourability (and to avoid self-loops in the graph), we must draw a path of edges between vertices in $S_{G}$ ($S_{B}$), where the path passes through non-green (non-blue) then green (blue) edges in an alternating fashion. We will add extra edges to the non-green (non-blue) parts of this path, while the green (blue) edges in this path will need to be removed in order to maintain tri-valence. For example, in Fig. 12(a), the dark green path connects vertices in $S_{G}$ and the light blue path connects vertices in $S_{B}$. The vertices in $S_{G}$ are not adjacent to a green edge, and hence we must add in green edges at these vertices, shown in Fig. 12(b). But this necessitates removing a green edge, and so on. The paths drawn in this figure are of length three: they each pass through two added edges and one removed edge. Finding the shortest such path minimises the number of removed connections.
Now consider how the algorithm from Sec. 3.1 forms such a path: it constructs it from the edges around a shrink plaquette. The two vertices in $S_{G}$ $(S_{B})$ will be separated by 3 edges that pass through the removed qubits. Hence, the length of the path constructed by the algorithm will be $K-3$ for a shrink plaquette of weight $K$. Hence the number of removed edges along the path will be $(K-4)/2$. In the honeycomb lattice, both shrink plaquettes will be weight 6, so the total number of additional removed edges (on top of the five adjacent to a removed qubit pair) will be 2. In the 4.8.8 lattice we also achieve 2 additional removed edges by choosing a defect edge such that one shrink plaquette has weight $K_{1}=8$ and another has weight $K_{2}=4$ (see Figs. 13(b) and 13(c) in Appendix C). In the 4.6.12 lattice we can achieve only 1 additional removed edge ($K_{1}=6$, $K_{2}=4$). In all cases it is obvious that there is no shorter path of edges (that does not pass through the removed qubits) that connects the vertices in $S_{G}$ and similarly for $S_{B}$. For example, in Fig. 12(a), a more optimal path would involve a direct connection between a pair of valence-2 vertices: no such edge can be formed. In the 4.8.8 case, a shorter path would imply a length-6 cycle of edges, which does not exist in the original lattice (assuming it is large enough that the non-trivial cycles are longer than six edges). ∎
In the above proof we have focused on uniform tilings on the plane, however this work can apply more generally to other surfaces as well. Below we remark how this optimality can also be shown for the 8.8.8 lattice, which is used in hyperbolic Floquet codes [48, 49].
Remark 1.
Consider a defective qubit in the bulk of a Floquet code defined on an 8.8.8 lattice (without boundary), which is a particular uniform tiling of a hyperbolic surface. Under the constraint that we cannot form an edge between any two qubits not connected by an edge in the original lattice, the algorithm described in Sec. 3.1 removes the defective qubit with the minimal number of removed edges from the original lattice.
Proof.
This proof follows from the fact that the proof for Theorem 2 does not rely on the lattice being on a plane. In the 8.8.8 lattice, the two shrink plaquettes will always be of weight-8, meaning that there will be 5 vertices in the paths connecting the vertices in $S_{G}$ and $S_{B}$, and a total of 4 edges are removed from the lattice (in addition to the five incident to the removed qubit pair). A more optimal path in the 8.8.8 lattice would require two degree-2 vertices being connected by a path shorter than length 5 (that does not pass through the removed qubits). This is not possible under the constraint that additional edges between non-connected vertices cannot be formed. ∎
@@
Hence, our scheme is also optimal in terms of the connections that are removed from the lattice. Note that we do not claim optimality for non-trivial combinations of defects, since then one must, for example, carefully choose the defect edges to achieve the optimal result. We do not have an optimal algorithm for such a choice, rather we use simplifying heursitics towards optimality (see Appendix D). Moreover, it may be true that there exists a path between vertices in $S_{G}$ or $S_{B}$ that is shorter than that constructed from the corresponding shrink plaquette, if that shrink plaquette is very large-weight (see Fig. 14(c) in Appendix D). We expect that finding optimality for non-trivial combinations of defects would not greatly improve code performance for small values of the defect probability.
@@ Appendix C Applicability to other Floquet codes
In this appendix, we discuss how the techniques described in Sec. 3 can be applied to other 2-D Floquet codes. Many 2-D Floquet codes can be defined on a colour code lattice, including the 4.8.8 Floquet code, the hyperbolic Floquet codes and the CSS Floquet code [4, 48, 49, 25, 24]. These codes can be adapted to accommodate defective qubits by selecting a defect edge incident to the defective qubit, along with selecting the merge and shrink plaquettes in the same way as in Steps 2 and 3 as described in Sec. 3.1, with the rest of the super-plaquette algorithm following naturally. An example of applying this approach to the 4.8.8 Floquet code is given in Fig. 13. The only step that may change is Step 4, wherein we choose bases for the added edges. However, if one can make an identification of edge basis with edge colour at any given measurement sub-round (see, e.g., Ref. [24]), then this modification is simple. For example, in the CSS Floquet code, one measures operators along coloured edges in the cyclical pattern $R\rightarrow G\rightarrow B$ with the basis of the measurement operator alternating cyclically as $X\rightarrow Z$. Hence, if we use our algorithm to adapt this Floquet code, the measurement schedule remains unchanged since we can consistently colour the newly added edges in the lattice.